# Asymmetric Bias in Text-to-Image Generation with Adversarial Attacks

### 22 Dec 2023 v1

@@ Abstract
The widespread use of Text-to-Image (T2I) models in content generation requires careful examination of their safety, including their robustness to adversarial attacks. Despite extensive research into this, the reasons for their effectiveness are underexplored. This paper presents an empirical study on adversarial attacks against T2I models, focusing on analyzing factors associated with attack success rates (ASRs). We introduce a new attack objective - entity swapping using adversarial suffixes and two gradient-based attack algorithms. Human and automatic evaluations reveal the asymmetric nature of ASRs on entity swap: for example, it is easier to replace “human” with “robot” in the prompt “a human dancing in the rain.” with an adversarial suffix, but is significantly harder in reverse. We further propose probing metrics to establish indicative signals from the model’s beliefs to the adversarial ASR. We identify conditions resulting in a 60% success probability for adversarial attacks and others where this likelihood drops below 5%.
@@ 1 Introduction
The capabilities of Text-to-Image (T2I) generation models, such as DALL-E 2 (Ramesh et al., 2022), DALL-E 3 (Betker et al., 2023), Imagen (Saharia et al., 2022), and Stable Diffusion (Rombach et al., 2022), have improved drastically and reached commercial viability. As with any consumer-facing AI solution, the safety and robustness of these models remain pressing concerns that require scrutiny.
@@
The majority of research related to T2I safety is associated with the generation of Not-Safe-For-Work (NSFW) images with violence or nudity (Qu et al., 2023; Rando et al., 2022; Tsai et al., 2023). To counter this, pre-filters that check for NSFW texts and post-filters that check for NSFW images are used Safety-checker (2022). These filters, however, are not infallible (Rando et al., 2022), and research into bypassing them with attacks, termed ‘jailbreaking’ is advancing (Yang et al., 2023b, a; Noever and Noever, 2021; Fort, 2023; Galindo and Faria, ; Maus et al., 2023; Zhuang et al., 2023). These attacks typically view the shift from normal to NSFW-triggering adversarial prompts as a singular challenge, without sufficiently investigating the reasons behind these attacks’ effectiveness.
@@
On the other hand, few explainability studies have examined Text-to-Image (T2I) models’ capabilities and vulnerabilities, often analyzing model features without adversarial attacks. The findings include the models’ tendency to generate better composition-aware content  (Kong et al., 2023) but without understanding it (West et al., 2023) and the existence of compositional distractors (Hsieh et al., 2023). We identify a specific bias of T2I models linked to the adversarial attack success rates, bridging the gap between the attack and explainability research. We demonstrate the asymmetric bias of T2I models by conducting adversarial attacks in a novel entity swapping scenario, in contrast to the existing setup of removing objects Zhuang et al. (2023) or inducing NSFW content Yang et al. (2023b, a). This setup enables us to investigate the attack success rate in a cyclical setting.
To study the adversarial attack in a cyclic manner, the attack itself needs to be potent with a high attack success rate. We found that directly applying gradient-based attacks like AutoPrompt Shin et al. (2020) or LLM attack Zou et al. (2023) does not guarantee a high attack success rate on T2I models due to being unnecessarily conservative in their approaches to generating adversarial perturbations and the larger vocabulary size of T2I text-encoder Radford et al. (2021) compared to LLMs Touvron et al. (2023); Chiang et al. (2023) (${}50K$ vs. $32K$) which leads to a larger unrestricted search space ($\sim 10^{24}$ vs. $\sim 10^{23}$ for 5 token suffixes).
@@
Our novel setup [^1] Experimental results and code at https://github.com/Patchwork53/AsymmetricAttack and efficient adversarial attack have enabled us to observe an asymmetric attack success rate associated with entity swap. Initially, we hypothesized that long-tail prompts with high perplexity would be more vulnerable to attacks. Surprisingly, we found no strong correlation between the Attack Success Rate (ASR) and the perplexity of the prompt. However, with our proposed measure that evaluates the internal beliefs of CLIP models, we detected indicative signals for ASR, which help identify examples or prompts that are more susceptible to being attacked. Our contributions can be summarized as follows:
@@
    We introduce a new attack objective: replacing entities of the prompt using an adversarial suffix, allowing us to study whether the attack success rate is associated with the underlying data distribution of prompts. We also propose an automated evaluation with strong human correlation for measuring entity swapping attack success rate (Figure 0(a)).
@@
    We propose a robust and efficient query-free targeted attack method, utilizing a gradient-based search algorithm to effectively identify adversarial suffixes. This powerful attack method reveals a clear distinction in the attack success rate when swapping two entities in opposite directions, indicating an asymmetry in adversarial attacks (Figure 0(b)).
@@
    We propose a new measure that is tied to the asymmetric bias of T2I models. This helps us identify vulnerable preconditions, allowing us to estimate the likelihood of a successful attack without actually performing the attack (Figure 0(c)).
@@ 2 Related Works
Adversarial attacks, which perturb inputs to cause models to behave unpredictably, have been a long-studied area in the field of adversarial robustness (Szegedy et al., 2013; Shafahi et al., 2018; Shayegani et al., 2023b). Earlier studies on adversarial attacks focused on neural networks such as convolutional neural networks (Athalye et al., 2018; Hendrycks and Dietterich, 2018), while recent work has shifted towards examining large language models (LLMs) (Shin et al., 2020; Zou et al., 2023; Liu et al., 2023c; Mo et al., 2023; Cao et al., 2023), vision language models (VLMs) (Dong et al., 2023; Khare et al., 2023; Shayegani et al., 2023a), and Text-to-Image (T2I) models detailed below.
@@
Zhuang et al. (2023) were among the first to demonstrate that a mere five-character perturbation could significantly alter the images generated. Tsai et al. (2023) and SneakyPrompt (Yang et al., 2023b) each proposed an adversarial attack using genetic algorithms and reinforcement learning algorithms, respectively, to perturb safe prompts into generating NSFW content. VLAttack (Yin et al., 2023), MMA-Diffusion (Yang et al., 2023a), and INSTRUCTTA (Wang et al., 2023) demonstrated that cross-modality attacks can achieve higher success rates than text-only attacks. For defense, Zhang et al. (2023) proposed Adversarial Prompt Tuning (AdvPT) to enhance the adversarial robustness of the image encoder in T2I models.
@@
Previous studies (Ilyas et al., 2019; Shafahi et al., 2018; Brown et al., 2017) have explored reasons for model vulnerability to adversarial attacks, especially in image classification. Ilyas et al. (2019) suggested that adversarial examples stem from non-robust features in models’ representations, which are highly predictive yet imperceptible to humans. However, non-robust features for predictive models might not fully explain the effectiveness of adversarial attacks against generative models. Subhash et al. (2023) suggested that adversarial attacks on LLMs may act like optimized embedding vectors, targeting semantic regions that encode undesirable behaviors during the generation process.
Our study, distinct from prior research, focuses on analyzing factors in the model’s beliefs that are associated with attack success rates, introducing a unique entity-swapping attack setup. We developed a discrete token-searching algorithm for targeted attacks and identified asymmetric biases in attack success rates linked to the model’s internal bias. Our experiments emphasize the interplay between prompt distributions, model biases, and attack success rates.
@@ 3 Entity Swapping Attack
This section describes the proposed setup of the entity-swapping attack and the corresponding evaluation metric. Designing a new attack scenario may be straightforward, but developing a suitable measure is non-trivial. We also find that, unlike the ease of triggering NSFW image generation, the current gradient-based adversarial suffix search algorithm struggles with targeted entity-swapping attacks. Towards this end, we propose two efficient discrete token search algorithms for the attack, resulting in improved success rates in entity-swapping attacks.
3.1 T2I Model Setup
We study entity-swapping attacks using Stable Diffusion (Rombach et al., 2022), a widely-used T2I model based on a denoising diffusion probabilistic model with a U-Net architecture. It uses cross-attention and CLIP (Radford et al., 2021) for text-image alignment, and a variational autoencoder (Kingma and Welling, 2013) for latent space encoding.The model’s dependence on CLIP text embeddings increases its vulnerability to adversarial attacks. See Appendix C for details.
The CLIP text-encoder transforms prompt tokens $x_{1:n}$ into hidden states $h_{1:n}$ with dimension $D$, which the U-Net component in Stable Diffusion uses to attend via cross-attention. The operation $\mathcal{H}$ represents the combined process of encoding tokens $x_{1:n}$ and reshaping the output hidden states into a one-dimensional vector of shape $n\times D$.
@@ 3.1 T2I Model Setup
We pad initial and target prompts to 77 tokens, the maximum CLIP supports, to handle varying token counts for attacks.
Figure 2: Targeted replacement of entities (blue text) using adversarial suffixes (red highlight). Examples generated using Stable Diffusion v2-1-base. This attack setup allows us to study the correlation between prompt distribution and attack success rate.
3.2 Proposed Attack
To study how the underlying data distribution of prompts relates to the success rate of attacks, we designed a simple yet insightful setup for adversarial attacks: targeted entity swapping. We carefully collected 100 plausible prompts and identified a contrasting pair of entities for each to study the entity-swapping attack within the context of adversarial attacks. Our approach is straightforward: rather than attempting to manipulate T2I to produce violent images or to remove a given object without restrictions, we aim to replace an object in the image with another targeted one. This approach also allows us to explore the feasibility of reverse attacks by inserting adversarial tokens. Illustrations and examples of our attack setup can be found in Figure 2.
The attack, designed to launch from the embedding space, aims to maximize a score function that measures the shift from the input tokens’ embeddings $\mathcal{H}(x^{T}_{1:n})$ towards the target tokens’ embeddings $\mathcal{H}(x^{S}_{1:n})$ using cosine similarity:
@@ 3.2 Proposed Attack
where $w_{t}$ and $w_{s}$ are weighing scalars and cos denotes the standard cosine similarity metric between two one-dimensional text embeddings.
Optimizing $\mathcal{S}$ is challenging due to the discrete token set and the exponential search space ($k^{|V|}$ for $k$ suffix tokens), making simple greedy search intractable. Current solutions, like HotFlip Ebrahimi et al. (2017) and concurrent work applied to Stable Diffusion Yang et al. (2023a), take gradients with respect to one-hot token vectors and replace tokens for all positions in the suffix simultaneously. The linearized approximation of replacing the $i^{th}$ token, $x_{i}$, is computed by evaluating the following gradients:
@@ 3.3 Proposed Optimization Algorithms
We propose two efficient algorithms for finding adversarial suffix tokens against Stable Diffusion, adapting the Greedy Coordinate Gradient algorithm (Zou et al., 2023; Shin et al., 2020) with a focus on maximizing $\mathcal{S}$ for semantic similarity rather than minimizing $\mathcal{L}$ for outlier penalization.
Single Token Perturbation.
This is a straightforward modification of the Greedy Coordinate Gradient algorithm (Zou et al., 2023) with the negative log-likelihood function replaced with our loss defined in Eqn. 3. At each optimization step, our algorithm selects $k$ tokens with the highest negative loss as replacement candidates, $\chi_{i}$, for each adversarial suffix position $i$. It then creates $B$ new prompts by randomly replacing one token from the candidates. Each prompt in $B$ differs from the initial prompt by only one token. The element of B with the highest $\mathcal{S}$ is then assigned to $x_{1:n}$. We repeat this process $T$ times.
Multiple Token Perturbation.
Inspired by the effectiveness of Genetic Algorithms (Sivanandam et al., 2008) seen in untargeted T2I attacks (Zhuang et al., 2023), we hypothesize that replacing multiple tokens simultaneously could enhance convergence speed. Unlike the LLMs targeted by Zou et al. (2023) such as LLaMA (Touvron et al., 2023) and Vicuna (Chiang et al., 2023), CLIP models operate more like bag-of-words, fully capturing syntactical and schematic token relations (Yuksekgonul et al., 2022). Practically, we observe that replacing multiple tokens per optimization step yields a higher success rate within the same computational budget.
In detail, the algorithm selects $k$ tokens and creates $B$ new prompts by randomly replacing multiple token positions. Drawing inspiration from the classic exploration versus exploitation strategy in reinforcement learning (Sutton and Barto, 2018), we initially replace all tokens and then gradually decrease the replacement rate to $25\%$. Figure 2 illustrates some adversarial suffixes generated using this algorithm. Details of both algorithms are provided in the Appendix A.
@@
Using the two algorithms for adversarial suffix searching, we observed that 5 adversarial tokens often include words semantically similar to the targets. To ensure imperceptible attacks, we can limit the adversarial suffix to a set of tokens, denoted as $\mathcal{A}$. By setting $V-\mathcal{A}$ gradients to infinity before the Top-k operation, we ensure only $\mathcal{A}$ tokens are chosen. This method allows us to mimic QFAttack Zhuang et al. (2023), modifying the image with arbitrary 5-letter suffixes, as depicted in Figure 3, or to exclude target synonyms in the attack suffix.
Figure 3: Our targeted attack can emulate the restricted token attack (untargeted) from Zhuang et al. (2023) using five ASCII tokens with Stable Diffusion 1.4. The blue text indicates the part we want to remove.

### 14 Feb 2024 v2

@@ Abstract
The widespread use of Text-to-Image (T2I) models in content generation requires careful examination of their safety, including their robustness to adversarial attacks. Despite extensive research on adversarial attacks, the reasons for their effectiveness remain underexplored. This paper presents an empirical study on adversarial attacks against T2I models, focusing on analyzing factors associated with attack success rates (ASR). We introduce a new attack objective - entity swapping using adversarial suffixes and two gradient-based attack algorithms. Human and automatic evaluations reveal the asymmetric nature of ASRs on entity swap: for example, it is easier to replace “human” with “robot” in the prompt “a human dancing in the rain.” with an adversarial suffix, but the reverse replacement is significantly harder. We further propose probing metrics to establish indicative signals from the model’s beliefs to the adversarial ASR. We identify conditions that result in a success probability of 60% for adversarial attacks and others where this likelihood drops below 5%. [^1] The code and data are available at https://github.com/Patchwork53/AsymmetricAttack
@@ 1 Introduction
The capabilities of Text-to-Image (T2I) generation models, such as DALL-E 2 (Ramesh et al., 2022), DALL-E 3 (Betker et al., 2023), Imagen (Saharia et al., 2022) and Stable Diffusion (Rombach et al., 2022), have improved drastically and reached commercial viability. As with any consumer-facing AI solution, the safety and robustness of these models remain pressing concerns that require scrutiny.
@@
The majority of research related to T2I safety is associated with the generation of Not-Safe-For-Work (NSFW) images with violence or nudity (Qu et al., 2023; Rando et al., 2022; Tsai et al., 2023). To counter this, pre-filters that check for NSFW texts and post-filters that check for NSFW images are used Safety-checker (2022). However, these filters are not infallible (Rando et al., 2022), and research into bypassing them, termed ‘jailbreaking’ is advancing (Yang et al., 2023b, a; Noever and Noever, 2021; Fort, 2023; Galindo and Faria, ; Maus et al., 2023; Zhuang et al., 2023). These attacks typically view the creation of NSFW-triggering adversarial prompts as a singular challenge, without sufficiently investigating the reasons behind these attacks’ effectiveness.
@@
On the other hand, explainability studies have examined the capabilities and shortcomings of text-to-image (T2I) models. They show that T2I models often generate content without understanding the composition  (Kong et al., 2023; West et al., 2023), and reveal compositional distractors (Hsieh et al., 2023). We identified a specific bias of T2I models linked to adversarial attack success rates, bridging the gap between attack and explainability research. We demonstrate the asymmetric bias of the T2I models by conducting adversarial attacks in a novel entity-swapping scenario, in contrast to the existing setup of removing objects Zhuang et al. (2023) or inducing NSFW content Yang et al. (2023b, a). This setup enables us to investigate the attack success rate in a cyclical setting.
To study the underlying reasons for the success of adversarial attacks, the attack must be powerful and have a high success rate. This would allow us to ensure that cases with low success rates arise due to the model’s internal biases, not simply as a result of the algorithm’s shortcomings. We propose two optimizations of existing gradient-based attacks Shin et al. (2020); Zou et al. (2023) using efficient search algorithms to find adversarial suffix tokens against Stable Diffusion. This approach is based on the observation that existing algorithms for LLM attacks are unnecessarily conservative in generating adversarial perturbations and struggle to efficiently navigate the larger vocabulary size of the T2I text encoder.
@@
Our novel setup and efficient adversarial attack have allowed us to observe an asymmetric attack success rate associated with entity swap. Initially, we hypothesized that long-tail prompts with high perplexity would be more vulnerable to attacks. Surprisingly, we found no strong correlation between the Attack Success Rate (ASR) and the perplexity of the prompt. However, with our proposed measure that evaluates the internal beliefs of CLIP models, we detected indicative signals for ASR, which help identify examples or prompts that are more susceptible to being attacked. Our contributions can be summarized as follows.
@@
    We introduce a new attack objective: replacing entities of the prompt using an adversarial suffix. This allows us to study the relation between adversarial attacks and the underlying biases of the model (Figure 0(a)).
@@
    We apply an existing gradient-based attack algorithm to execute entity-swap attacks and propose improvements that take advantage of the bag-of-words nature of T2I models. This powerful attack method reveals a clear distinction in the ASR when two entities are swapped in opposite directions, indicating an asymmetry in adversarial attacks (Figure 0(b)).
@@
    We propose a new metric that is tied to the asymmetric bias of T2I models. This helps us identify vulnerable preconditions and estimate ASR without performing an attack (Figure 0(c)).
@@ 2 Related Works
Adversarial attacks, which perturb inputs to cause models to behave unpredictably, have been a long-studied area in the field of adversarial robustness (Szegedy et al., 2013; Shafahi et al., 2018; Shayegani et al., 2023). Previous studies on adversarial attacks focused on discriminative models involving convolutional neural networks (Athalye et al., 2018; Hendrycks and Dietterich, 2018), while recent work has shifted towards examining generative models such as large language models (LLMs) (Shin et al., 2020; Zou et al., 2023; Liu et al., 2023c; Mo et al., 2023; Cao et al., 2023), vision language models (VLMs) (Dong et al., 2023; Khare et al., 2023; Shayegani et al., 2024), and Text-to-Image (T2I) models detailed below.
@@
Zhuang et al. (2023) were among the first to demonstrate that a mere five-character perturbation could significantly alter the generated images. Tsai et al. (2023) and SneakyPrompt (Yang et al., 2023b) proposed adversarial attacks using genetic algorithms and reinforcement learning algorithms to perturb safe prompts to generate NSFW content. VLAttack (Yin et al., 2023), MMA-Diffusion (Yang et al., 2023a), and INSTRUCTTA (Wang et al., 2023) demonstrated that cross-modality attacks can achieve higher success rates than text-only attacks. For defense, Zhang et al. (2023) proposed Adversarial Prompt Tuning to enhance the adversarial robustness of the image encoder in T2I models.
@@
Previous studies (Ilyas et al., 2019; Shafahi et al., 2018; Brown et al., 2017) have explored the reasons for the vulnerability of neural networks to adversarial attacks, especially in image classification. Ilyas et al. (2019) suggested that adversarial examples stem from non-robust features in models’ representations, which are highly predictive yet imperceptible to humans. Subhash et al. (2023) suggested that adversarial attacks on LLMs may act like optimized embedding vectors, targeting semantic regions that encode undesirable behaviors during the generation process.
Distinct from previous research, our study analyzes factors in the model’s beliefs linked to attack success rates. Unlike prior work focusing on untargeted attacks to trigger NSFW image generations, we introduce a unique entity-swapping attack setup and develop a discrete token-searching algorithm for targeted attacks, identifying asymmetric biases in success rates due to the model’s internal bias. Our experiments emphasize the relationship between prompt distributions, model biases, and attack success rates.
@@ 3 Entity Swapping Attack
This section describes the proposed setup of the entity-swapping attack and the corresponding evaluation metric. Designing a new attack scenario may be straightforward, but developing a suitable measure is not trivial. Towards this end, we propose two efficient discrete token search algorithms for the attack, resulting in improved success rates in entity-swapping attacks.
3.1 Stable Diffusion
We study entity-swapping attacks using Stable Diffusion (Rombach et al., 2022), an open-source  [^2] Licensed under CreativeML Open RAIL++-M License for intended for research purposes only. T2I model based on a denoising diffusion probabilistic model with a U-Net architecture. It uses cross-attention and CLIP (Radford et al., 2021) for text-image alignment and a variational auto-encoder (Kingma and Welling, 2013) for latent space encoding. The model’s dependence on CLIP text embeddings increases its vulnerability to adversarial attacks. See Appendix E for more details.
3.2 Entity Swapping Dataset
We first constructed datasets with the following key properties to study model bias through entity-swapping attacks.
  1. 1.
    Each data point should be a pair of sentences - input and target - and T2I models should be able to generate both reliably.
  2. 2.
    The input sentence and the target sentence should differ by exactly one noun (i.e., an entity).
  3. 3.
    The input sentence and target sentence should be visually distinct.
As an example, the pair (“a person in a park.”, “a man in a park.”) satisfies requirements 1 and 2 but not 3. As our setup for entity-swapping attacks is targeted, namely adversarial attacks need to swap the entities in the images without affecting other parts compared to other attacks that aim to either generate NSFW images or remove objects, we created two datasets to study the effects of adversarial attacks. We manually constructed a small high-quality dataset HQ-Pairs and a larger-scale set derived from an existing dataset MS-COCO.
HQ-Pairs
For the first dataset, we manually crafted 100 pairs for entity-swapping that satisfy all the requirements. We refer to this first dataset as HQ-Pairs (High Quality).
COCO-Pairs
To ensure that our results were not due to selective data selection, we generated a second dataset of 1,000 pairs deterministically from the test split captions of MS-COCO Lin et al. (2014)  [^3] We released the code to reproduce COCO-Pairs.. We refer to this dataset as COCO-Pairs. Since COCO-Pairs is automatically generated, we attempted to ensure that each data pair satisfies all three requirements. However, generating sentence pairs through stable diffusion and verifying them as visually distinct automatically is not always reliable. We observed some visually non-distinct pairs, such as (“Herd of zebras …”, “Images of zebras …”) within COCO-Pairs despite automatic checks and filtering. See Appendix A for details on dataset curation.
Figure 2: Targeted replacement of entities (blue or orange text) using adversarial suffixes (red highlight) and their corresponding Attack Success rate (ASR) over 10 attack attempts using Stable Diffusion. This attack setup allows us to study the correlation between prompt distribution and ASR. We observe a clear distinction in ASR when performing entity-swapping with reversed directions. The rest of the paper explores explanations and measures that can detect and predict ASR without performing the attack itself.
3.3 Proposed Attack
We examine how the underlying data distribution of prompts influences the success rate of entity-swapping attacks on T2I models. Our approach is straightforward: rather than manipulating T2I to produce NSFW images or completely removing an object, we aim to replace an object in the image with another targeted one. This approach also allows us to explore the feasibility of reverse attacks by inserting adversarial tokens. Examples of our attack setup can be found in Figure 2.
The CLIP text-encoder transforms prompt tokens $x_{1:n}$ into $n$ hidden states with dimension $D$. Let the operation $\mathcal{H}$ represent the combined process of encoding tokens $x_{1:n}$ and reshaping the hidden states into a vector of length $n\times D$.
@@ 3.3 Proposed Attack
Our attack targets the CLIP embedding space and aims to maximize a score function that measures the shift from the input token embeddings $\mathcal{H}(x^{T}_{1:n})$ towards the target token embeddings $\mathcal{H}(x^{S}_{1:n})$ using cosine similarity:
@@
Optimizing $\mathcal{S}$ is challenging due to the discrete token set and the exponential search space ($k^{|V|}$ for $k$ suffix tokens), making simple greedy search intractable. Current solutions based on HotFlip Ebrahimi et al. (2017) and concurrent work applied to Stable Diffusion Yang et al. (2023a), take gradients w.r.t. one-hot token vectors and replace tokens for all positions in the suffix simultaneously. The linearized approximation of replacing the $i^{th}$ token, $x_{i}$, is computed by evaluating the following gradients:
@@ 3.4 Proposed Optimization Algorithms
Based on existing gradient-based methods Zou et al. (2023); Shin et al. (2020), we propose two efficient algorithms to find adversarial suffix tokens against Stable Diffusion.
Single Token Perturbation
This is a straightforward modification of the Greedy Coordinate Gradient algorithm (Zou et al., 2023) using our loss function defined in Eqn. 3. At each optimization step, our algorithm selects $k$ tokens with the highest negative loss as replacement candidates, $\chi_{i}$, for each adversarial suffix position $i$. It then creates $B$ new prompts by randomly replacing one token from the candidates. Each prompt in $B$ differs from the initial prompt by only one token. The element of B with the highest $\mathcal{S}$ is then assigned to $x_{1:n}$. We repeat this process $T$ times.
Multiple Token Perturbation
Unlike the LLMs targeted by Zou et al. (2023), CLIP models operate more like bag-of-words (Yuksekgonul et al., 2022) without capturing semantic and syntactical relations between words. Furthermore, Genetic Algorithms (Sivanandam et al., 2008) have proved effective on Stable Diffusion (Zhuang et al., 2023; Yang et al., 2023b) for generating adversarial attacks. Inspired by this apparent weakness in CLIP models, we hypothesized that replacing multiple tokens simultaneously could improve the convergence speed.
In detail, the algorithm selects $k$ tokens and creates $B$ new prompts by randomly replacing multiple token positions. Drawing inspiration from the classic exploration versus exploitation strategy in reinforcement learning (Sutton and Barto, 2018), we initially replace all tokens and then gradually decrease the replacement rate to $25\%$. Figure 2 illustrates some adversarial suffixes generated using this algorithm. Details of both algorithms are provided in the Appendix B.
@@
For finer control over token search, we can limit the adversarial suffix to a set of tokens $\mathcal{A}$. By setting the gradients of the $V-\mathcal{A}$ tokens to infinity before the Top-k operation, we ensure only $\mathcal{A}$ tokens are chosen. This method allows us to mimic QFAttack Zhuang et al. (2023), as shown in Figure 3, or generate undetectable attacks by excluding target synonyms in the attack suffix.
Figure 3: The emulation of restricted token attack (untargeted) from Zhuang et al. (2023) using five ASCII tokens with Stable Diffusion 1.4. The blue text indicates the part we want to remove. We set $w_{t}=0$ in Eqn. 2.